VIS (IX)takaaki.yokozawa - 10:41 Wednesday 21 May 2025 (33848)
Print this reportMorning ITMX checkThis is a log of work in this morning.
I requested the ISOLATED and performed the TM TILT centering using the optic align. With the P=4300cnt, Y=-1820, TM TILT QPD became almost center. After centering, I performed the OLG check for the MN DAMP filter P. As shown in Fig.1., the OLG shape seemed much better and 7.5 Hz peak can be seen. (I didn't say it became healthy)
After that I performed the offload work using the moving mass and BF yaw setting point.
Images attached to this report
Comments to this report:
takaaki.yokozawa - 10:57 Wednesday 21 May 2025 (33851)
Print this reportAfter two hours, the OLG shape changed very much. So, close to health, but not health now. Tanaka-san will report the result of health check later.